OpenStudy (anonymous):

I just need help setting this problem up: Find the double integral by interpreting it as the volume of a solid.

OpenStudy (anonymous):

\[\int\limits_{}^{}\int\limits_{}^{}3-\frac{ 1 }{ 2 }x-\frac{ 3 }{ 4 }y \]

OpenStudy (anonymous):

R= 2x+3y is less than or equal to 12 and both x and y are greater than 0

OpenStudy (anonymous):

What I tried doing is this: \[\int\limits_{0}^{6}\int\limits_{0}^{-2/3x+4}f(x,y) dy dx\]

OpenStudy (anonymous):

but I got a negative answer

zepdrix (zepdrix):

Hmmmm seems like you're setting it up correctly.

zepdrix (zepdrix):

And the function is the thing you listed first, yes?

OpenStudy (anonymous):

yes, it is. I guess I'm glad I'm setting it up right :/

OpenStudy (anonymous):

as long as I know that, I'll just keep reworking it. Thank you :)

OpenStudy (anonymous):

I may try doing dx first too

zepdrix (zepdrix):

Just make sure your outer integral contains the `constant` boundaries. Yah doing it in dx might give you some more perspective, just make sure your line is rewritten in terms of y.

OpenStudy (anonymous):

definitely. Thank you. I was afraid I wasn't understanding the concept.